OpenStudy (anonymous):

Why can’t we use the intercept method to graph the following equation? 3(x+2) = 2(y+3)

OpenStudy (anonymous):

@Hero

hero (hero):

Have you attempted to use the intercept method to graph the equation yet?

OpenStudy (anonymous):

A bit but i cant really understand

hero (hero):

The intercept method works like this: To find the x intercept, let y = 0, then solve for x. The point you find will be (x, 0) where x is the value you solved for. To find the y intercept, let x = 0, then solve for y. The point you find will be (0, y) where y is the value you solved for. Each point is usually two distinct points as you only need two points two draw a line. For example: Suppose you had to graph y = 2x + 4 Finding the x intercept, you let y = 0, the solve for x as follows: 0 = 2x + 4 -4 = 2x -4/2 = x -2 = x So the x intercept is the point (-2, 0) To find the y intercept, let x = 0, then solve for y as follows: y = 2(0) + 4 y = 0 + 4 y = 4 So the y-intercept is the point (0 , 4)

hero (hero):

You would then proceed to plot those two distinct points, then draw a line through them and label the line y = 2x + 4

hero (hero):

In this case, for the equation you posted, if you tried to find two distinct intercepts, you would realize that the x-intercept = y-intercept = (0,0).

OpenStudy (anonymous):

ok i understand but y couldn't i do that same thing with 3(x+2) = 2(y+3)

hero (hero):

In other words, you only have one point rather than the required two points you need to draw a line. Remember two points gives you a slope which is a necessary property of a line. With only one point, you won't know the slope of the line, therefore there would be no way to draw a line.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks